OpenStudy (anonymous):

The shorter leg of a 30-60-90 triangle is 6. What is the length of the other leg?

OpenStudy (anonymous):

30(a-5) a=5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

short leg is 6, hypotenuse is 12 and other leg is \[\sqrt{12^2-6^2}\]

OpenStudy (anonymous):

i get \[\sqrt{108}\]

OpenStudy (anonymous):

maybe im not sure!!! :(

OpenStudy (anonymous):

otherwise known as \[6\sqrt{3}\]

OpenStudy (anonymous):

my ans. is 5?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the answer is \[6\sqrt{3}\]

OpenStudy (anonymous):

if you are using trig you know that \[tan(30)=\frac{6}{x}=\frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

making \[x=6\sqrt{3}\]

OpenStudy (anonymous):

ratios of 30-60-90 triangle are hypotenuse is 1, short side is 1/2 and long side is \[\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

here your short side is , so long side is \[6\sqrt{3}\]